OpenStudy (anonymous):

which point does the line for the equation y=x+2 pass through if you help me ill give you the graph thx

OpenStudy (jozelynw):

ok i will help

OpenStudy (jozelynw):

where is the pics

OpenStudy (anonymous):

aww thx

OpenStudy (anonymous):

let me attach it hold on

OpenStudy (jozelynw):

k

OpenStudy (jozelynw):

i wont leave like the other people do

OpenStudy (jozelynw):

Ok, I remember this I believe hold on let me look at my notes.

OpenStudy (anonymous):

okay

OpenStudy (jozelynw):

let me ask is this parallel and perpendicular lines

OpenStudy (anonymous):

no it is a graph

OpenStudy (jozelynw):

are you on flvs

OpenStudy (jozelynw):

what lesson is this on?

OpenStudy (anonymous):

yes im in connexus flvs

OpenStudy (jozelynw):

lesson

OpenStudy (anonymous):

unit 9 lesson 10

OpenStudy (jozelynw):

okay i cant help you because you are not in the same flvs as me and i do not understand this question but someone who will help @mathmate @peachpi good luck sorry.

OpenStudy (anonymous):

The first number is the x-coordinate and the second is the y.

OpenStudy (jozelynw):

she is only in 5th @peachpi

OpenStudy (mathmate):

A point (x,y) is on the line if the x and y satisfy the given equation y=mx+b For example, we have a line y=2x+1 then (0,0) is not on the line because 0\(\ne\)2(0)+1, or 0\(\ne\)1 But (1,3) is on the line because 3=2(1)+1 All you need to do is to subsiitute the x, y values into the given equation. Hope that helps.

OpenStudy (jozelynw):

grade

OpenStudy (anonymous):

so whats the awnser?

OpenStudy (anonymous):

im only in 5th grade I don't understand this

OpenStudy (jozelynw):

ok what about this @peachpi and @mathmate when you substitute y for zero and do the equation like that what about that

OpenStudy (anonymous):

huh?

OpenStudy (mathmate):

I'll check the first point for you. for the point (0,3) means x=0, y=3 Calculate: y=3 x+2=0+2=2 Since 3\(\ne\)2, we conclude that (0,3) does not lie on the line y=x+2. You can repeat the calculation for the other three points, and decide.

OpenStudy (jozelynw):

I'm trying but you are only in the 5th and this question is not 5th grade level

OpenStudy (jozelynw):

have you learned any of what we are saying @prettypink_mega

OpenStudy (anonymous):

no idea sorry 😡

OpenStudy (anonymous):

can you guys just give me the awnser because my mom will kill me if I fail

OpenStudy (jozelynw):

We don't even know the answer . Ask your sister

OpenStudy (anonymous):

shes out with my parents

OpenStudy (jozelynw):

when i google the question it pops up as perpendicular and parallel lines

OpenStudy (mathmate):

@JozelynW is right. Grade 5 students don't do this kind of work. I have explained how you can do it. You don't even need a calculator.

OpenStudy (anonymous):

can you ask your parents please I really need to pass this exam its the last assighment for math

OpenStudy (jozelynw):

girl my parents stopped knowing how to do my math work when I hit 6th grade,

OpenStudy (jozelynw):

try @mathmate way

OpenStudy (jozelynw):

or let me try it

OpenStudy (jozelynw):

I'll check the first point for you. for the point (0,3) means x=0, y=3 Calculate: y=3 x+2=0+2=2 Since 3≠2, we conclude that (0,3) does not lie on the line y=x+2. You can repeat the calculation for the other three points, and decide.

OpenStudy (anonymous):

The x is the number running across on the graph. That's the first number in parentheses. The y is the number running up and down. That's the 2nd number. To plot the point (0,3) move over by 0 and up by 3. To plot (2, 0) move right 2 and up 0. Plot the other 2 points and the answer is the on that touches the line.